OpenStudy (anonymous):

The function f changes value when x changes from a to a + dx. Find the appoximation error. F(x) = x^3, a =8, dx = 0.05

OpenStudy (anonymous):

what does u mean by x^3 ??? x*3 or x/3

OpenStudy (anonymous):

F with exponent 3

OpenStudy (anonymous):

Do u know how to solve?

OpenStudy (anonymous):

u have to take the cube of x

OpenStudy (anonymous):

Yes what is x

OpenStudy (anonymous):

As u said above that the value of x changes from a to a+dx So according to the given value x will be 8.05

OpenStudy (anonymous):

Ok my options for the answer are as following

OpenStudy (anonymous):

0.12025

OpenStudy (anonymous):

0.0300625

OpenStudy (anonymous):

0.09010875

OpenStudy (anonymous):

Or 0.060125

OpenStudy (anonymous):

I think 0.12025

OpenStudy (anonymous):

I think I am answering something else your question is not, what I am thinking

OpenStudy (anonymous):

Are u sure, cuz this is actually on a test im taking

OpenStudy (anonymous):

You should wait for someone else to answer it

OpenStudy (anonymous):

Ha okay

OpenStudy (ybarrap):

The exact answer is to take the derivative: F'(x) = 3x^2 F'(8) = 3*8^2=3*64=192 So now let's approximate: (F(8+.05) - F(8)) / .05 = ((8+.05)^3 - 8^3) / .05 = 193.2025 So the error in this approximation method is: (192 - 193.2025)/192 = -0.0062630208333... This is a pretty small!